Financial Planning and Forecasting: Using Regression to Improve Forecasts Regression analysis is a statistical technique that fits a line to observed data points so that the resulting equation can be used to forecast other data points. It is useful in excess capacity and economies of scale situations. Economies of scale occur when the ratio of asset to sales will change as the size of the firm increases. Regression analysis can lead to improved financial forecasts and better information which can be used to improve management's actions. Quantitative Problem: Jasper Jewelry has $160 million in sales. The company expects that its sales will increase 4% this year. Jasper's CFO uses a simple linear regression to forecast the company's inventory level for a given level of projected sales. On the basis of recent history, the estimated relationship between inventories and sales (in millions of dollars) is as follows: Inventories = $11 + 0.09(Sales) Given the estimated sales forecast and the estimated relationship between inventories and sales, what is your forecast of the company's year-end inventory level? Round your answer to two decimal places. Do not round intermediate calculations. $ million
Answer #1

Sales for the present period: $160 million

Expected increase in sales in the next period: 4%

Sales in the next period = 160*1.04 = $166.4 million

Estimated relationship between inventories and sales (in millions of dollars):

Inventories = $11 + 0.09(Sales)

Inventory for present period = $11 + 0.09(160) = $25.4 million

Inventory for next period = $11 + 0.09(166.4) = $25.976 million

Answer: Forecast for company's year-end inventory level = $25.976 million